OpenStudy (anonymous):

write the slope-intercept equation for the line with slope -2/5 and y-intercept (0,-3) and graph the line

OpenStudy (anonymous):

y - (-3 ) = -2/5(x-0) y + 3 = -2/5x y = -2/5x - 3

OpenStudy (callisto):

slope-int equation: y= mx+c -(1) where m is the slope and c is the y-intercept So, put m=-2/5 and c=-3 into (1) you can get the answer directly~